Generation Y has been defined as those individuals who were born between 1981 and 1991 A 2010 survey by a credit counseling foundation ound that 55% of the young adults in Generation Y pay their monthly bills on time Suppose we take a random sample of 210 people from Generation Y. Complete parts a through e below. a. Calculate the standard error of the proportion. Ơp-L (Round to four decimal places as needed.) b. What is the probability that 130 or fewer will pay their monthly bills on time? P(130 cr fewer Generation Y individuals will pay their monthlybils on time) (Round to four decimal places as needed.) c. What is the probability that 105 or fewer will pay their monthly bills on time? P(105 or fewer Generation Y individuals ill payheir monthly bills on time) (Round to four decimal places as needed.) d. What is the probability that 123 cr more will pay their monthly bills on time? P(123 or more Generation Y individuals will pay their monthly bills on time)- Round to four decimal places as needed.) e. What is the probability that between 113 and 122 of them will pay their monthly bills on time? P(Between 113 and 122 of them will pay their monthly bills on time) (Round to four decimal places as needed.)
Answer #1

a)

std error of proportion=σp=√(p*(1-p)/n)= 0.0343

b)

probability = P(X<0.619) = P(Z<2.01)= 0.9778

c)

probability = P(X<0.5) = P(Z<-1.46)= 0.0721

d)

probability = P(X>0.586) = P(Z>1.04)= 1-P(Z<1.04)= 1-0.8509= 0.1492

e)

probability = P(0.538<X<0.581) = P(-0.35<Z<0.9)= 0.8159-0.3632= 0.4527